# Active Inference AI Systems for Scientific Discovery

### v2

@@ Abstract
The rapid evolution of artificial intelligence has led to expectations of transformative scientific discovery, yet current systems remain fundamentally limited by their operational architectures, brittle reasoning mechanisms, and their separation from experimental reality. Building on earlier work, this perspective contends that progress in AI-driven science now depends on closing three fundamental gaps—the abstraction gap, the reasoning gap, and the reality gap—rather than on model size/data/test time compute. Scientific reasoning demands internal representations that support simulation of actions and response, causal structures that distinguish correlation from mechanism, and continuous calibration. Active inference AI systems for scientific discovery are defined as those that (i) maintain long-lived research memories grounded in causal self-supervised foundation models, (ii) symbolic or neuro-symbolic planners equipped with Bayesian guardrails, (iii) grow persistent knowledge graphs where thinking generates novel conceptual nodes, reasoning establishes causal edges, and real-world interaction prunes false connections while strengthening verified pathways, and (iv) refine their internal representations through closed-loop interaction with both high-fidelity simulators and automated laboratories—an operational loop where mental simulation guides action and empirical surprise reshapes understanding. In essence, this work outlines an architecture in which discovery arises from the interplay between internal models that enable counterfactual reasoning and external validation that grounds hypotheses in reality. It is also argued that the inherent ambiguity in feedback from simulations and experiments, and underlying uncertainties makes human judgment indispensable, not as a temporary scaffold but as a permanent architectural component.
1 Present day AI Systems and Scientific Discovery
@@
Modern science also operates within a myriad of constraints that exist in economic, legal and social dimensions rather than physical laws. These constraints favor certain types of AI-driven discovery while effectively prohibiting others. Additionally, the structure of the modern scientific enterprise—with its emphasis on incremental, publishable units and citation metrics—may be fundamentally misaligned with the kind of patient, integrative thinking required for paradigm shifts. AI systems could theoretically ignore these social pressures, pursuing research programs too risky or long-term for humans. But this same freedom from social constraint raises the possibility of AI systems optimizing for discovery without broader goals accounted for. Perhaps more concerning, AI systems trained on existing scientific literature risk amplifying current biases and narrowing the space of explored ideas. In other words, AI systems might converge on well-studied paths [11, 40], reducing the rich variety of research directions to a handful of statistically probable avenues. Scientific progress demands not convergence but divergence—an explosion of hypotheses, methodologies, and frameworks that challenge orthodoxy. The challenge is in designing AI systems that expand rather than constrain the landscape of scientific imagination.
Against this backdrop, the remainder of this perspective piece is organized around three interlocking hurdles that scientific discovery architectures must clear: (i) the abstraction gap, which separates low-level statistical regularities from the mechanistic concepts on which scientists actually reason; (ii) the reasoning gap, which limits today’s models to correlation-driven pattern completion rather than causal, counterfactual inference; and (iii) the reality gap, which isolates computation from the empirical feedback loops that ultimately arbitrate truth. Each gap both constrains and amplifies the others: without rich abstractions there is little substrate for reasoning, and without tight coupling to reality even the most elegant abstractions may drift toward irrelevance.
2 The Abstraction Gap
@@
The gap between correlation and causation represents perhaps the most fundamental challenge in automated scientific discovery. While current models excel at finding statistical regularities, scientific understanding requires the ability to reason about interventions—to ask not just “what correlates with what?" but “what happens when we change this?"
Pearl’s causal hierarchy [48] distinguishes three levels of cognitive ability: association (seeing), intervention (doing), and counterfactuals (imagining). Current AI systems operate primarily at the associative level, occasionally reaching intervention through experimental design. True scientific reasoning requires all three, particularly the counterfactual ability to imagine alternative scenarios that violate observed correlations. This connects directly to ethologist Konrad Lorenz’s insight –first tied to learning systems by Scholkopf [57]– that reasoning is fundamentally about acting in imaginary spaces where we can violate the constraints of observed data. This mental experimentation—impossible in physical reality but trivial in the imagination—forms the basis of scientific law formation.
3 Unhobbling Intelligence
A certain level of consensus appears to be forming in the community that incremental scaling of present architectures may not deliver the qualitative leap that scientific discovery demands. Progress hinges on unhobbling—removing the design constraints that keep today’s models predictable, yet fundamentally limited—through concurrent advances in algorithms, speculation control, hardware co-design, and access models.
Algorithmic Gains
Future systems must balance the complementary modes of thinking and reasoning as first-class architectural principles. Thinking—or slow, iterative discovery of new patterns—demands (i) world-model agents that can explore counterfactual spaces through mental simulation [25]; (ii) curiosity-driven mechanisms that reward pattern novelty over immediate task performance; and (iii) patience parameters that prevent premature convergence. Reasoning—the fast, deterministic traversal of pattern graphs—demands (i) efficient knowledge graph architectures with learned traversal policies; (ii) neuro-symbolic stacks that maintain both continuous representations and discrete logical structures[38]; and (iii) caching mechanisms that transform expensive thinking outcomes into rapid reasoning primitives. The interplay between these modes mirrors how scientists alternate between exploratory experimentation (thinking) and theoretical derivation (reasoning).
@@ 3 Unhobbling Intelligence
Scientific progress thrives on disciplined risk: venturing beyond received wisdom while remaining falsifiable. Current alignment protocols deliberately dampen exploratory behaviour, biasing models toward safe completion of well-trodden trajectories. Controlled speculation frameworks—for example, curiosity-driven reinforcement learning [45] combined with Bayesian epistemic guards—could allow systems to seek novel hypotheses, flag them with calibrated uncertainty, and propose targeted experiments for arbitration. Mechanisms such as self-consistency voting [64], adversarial peer review, and tool-augmented chain-of-thought audits offer additional scaffolding to keep high-variance reasoning tethered to empirical reality.
Computational Inefficiency
@@
The von Neumann bottleneck—shuttling tensors between distant memory and compute—now dominates energy budgets [39]. Processing-in-memory fabrics [32], spiking neuromorphic cores that exploit event sparsity , analog photonic accelerators for low-latency matrix products, quantum samplers for combinatorial sub-routines [2] could open presently unreachable algorithmic spaces. Realising their potential outside of niche applications, however, will require co-design of hardware, software and algorithms and extensive community effort.
@@
4 Architecture for the Era of Experience
Empirical feedback complements formal reasoning by supplying information inaccessible to purely deductive systems, thereby expanding—rather than mechanically escaping—the set of testable scientific propositions. The interplay between formal systems and empirical validation creates a bootstrap mechanism that circumvents incompleteness and irreducibility constraints. This suggests that AI systems for discovery must be fundamentally open—not just to new data, but to surprise from reality itself. Scientific history abounds with internally coherent theories that later failed empirical tests, underscoring the indispensability of continuous validation against data. Current AI systems excel at interpolation within their training distributions but struggle with the extrapolation that defines discovery. This is exacerbated by the fact that many scientific domains are characterized by sparse, expensive data and imperfect simulators. Unlike language modeling where data is abundant, a single protein crystallography experiment might take months and cost thousands of dollars. Simulations help but introduce their own biases—the “sim-to-real gap" that plagues robotics extends to all of computational science. Our architecture must therefore implement a hybrid loop: physics priors guide ML surrogates, which direct active experiments, which update our understanding in continuous iteration.
Causal Models
The current paradigm of domain-specific foundation models—from protein language models to molecular transformers—represents significant progress in encoding domain knowledge. However, these models fundamentally learn correlational patterns rather than causal mechanisms. ChemBERTa can predict molecular properties through pattern matching but cannot simulate how modifying a functional group alters reaction pathways. AlphaFold predicts protein structures through evolutionary patterns but does not model the physical folding process.
Scientific discovery demands models that transcend pattern recognition to capture causal dynamics. A causal molecular model would not just recognize that certain molecular structures correlate with properties—it would explain how electron density distributions cause reactivity, and how thermodynamic gradients drive reactions. This causal understanding enables the counterfactual reasoning essential to science: predicting outcomes of novel interventions never seen in training data. This architectural choice has profound implications: foundation models scale with data and compute, but causal models scale with understanding. As we accumulate more structural data, foundation models improve at interpolation. As we refine causal mechanisms, foundation models improve at extrapolation—the essence of scientific discovery.
Physics priors
While generative models like Sora create visually compelling outputs, they lack physical consistency—objects appear and disappear, gravity works intermittently, and causality is merely suggested rather than enforced. Mitchell [42] states that without biases to prefer some generalizations over others, a learning system cannot make the inductive leap necessary to classify instances beyond those it has already seen. Such inductive biases or physics priors—can be built-in to ensure generated realizations obey conservation laws, maintain object permanence, and support counterfactual reasoning about physical interactions.
Active Inference AI Systems to Navigate Complex Scientific Questions
Many scientific phenomena exhibit chaotic dynamics, multiscale interactions, and emergent properties that defy reductionist analysis. Climate systems, biological networks, and turbulent flows operate across scales from molecular to planetary. Traditional ML approaches that assume smooth, well-behaved functions fail catastrophically in these domains. We need architectures that can reason across scales, identify emergent patterns, and know when deterministic prediction becomes impossible. No single formal or informal computational system can accomplish these tasks, and hence we propose an AI stack. An exemplar architecture is shown in Figure 1. Some of the components of the architecture include:
@@ 4 Architecture for the Era of Experience
The architectural principles outlined above find grounding in recent work on transformational scientific creativity. For instance, Schapiro et al. [55] formalize scientific conceptual spaces as directed acyclic graphs, where vertices represent generative rules and edges capture logical dependencies. This offers a concrete implementation pathway for the proposed dynamic knowledge graphs. Their distinction between modifying existing constraints versus fundamentally restructuring the space itself maps directly onto our architecture’s dual modes of reasoning (traversing established knowledge) and thinking (discovering new patterns that may violate existing assumptions). This convergence suggests that achieving transformational scientific discovery through AI systems requires systems capable of identifying and modifying the foundational axioms that constrain current scientific understanding—a capability our the active inference framework aims to provide through its stacked architecture and integration of models, empirical feedback, and human guidance.
It is acknowledged that while the AI system can– in principle– be operated autonomously through well-defined interfaces between components, human interaction and decisions can be expected to play a key role. The architectural principles outlined above find partial instantiation in contemporary systems, though none fully realize the complete vision of scientific intelligence. Appendix A examines some current implementations through the lens of our three-gap framework, and discusses both substantial progress and persistent limitations that illuminate the path forward. Appendix B gives high–level mathematical constructs for key components of the above system.
Challenges of Iterative Learning and Importance of Human Interactions
@@
5 Outlook
Active AI systems encompass external experience (empirical data) and internal experience (mental simulation). AI systems that can fluidly navigate between these modes will mark the transition from tools that find patterns to partners that discover principles. This perspective builds upon substantial progress in causal machine learning, active learning, and automated scientific discovery while addressing critical gaps. The causal machine learning community has made significant strides in developing methods for causal inference from observational data, with frameworks like Pearl’s causal hierarchy and recent advances in causal representation learning providing mathematical foundations for understanding interventions and counterfactuals. Similarly, active learning has evolved sophisticated strategies for optimal experimental design, while automated discovery systems have demonstrated success in specific domains such as materials science and drug discovery. However, these communities have largely operated in isolation, with causal methods focusing primarily on statistical inference rather than physical mechanism discovery, active learning optimizing for narrow uncertainty reduction rather than conceptual breakthrough, and automated discovery systems excelling at interpolation within known spaces rather than extrapolation to genuinely novel phenomena.
Current implementations prioritize task completion over understanding, optimization over exploration, and correlation over causation. The path forward requires AI systems that integrate causal reasoning not merely as a statistical tool but as the foundation for mental simulation and counterfactual experimentation, extending active learning beyond data efficiency to include the generation and testing of novel hypotheses that violate existing assumptions, and grounding automated discovery in continuous empirical feedback loops that prevent drift from physical reality. Most critically, while existing approaches excel within their prescribed domains, they lack the architectural foundation for the kind of open-ended, cross-domain reasoning that characterizes human scientific discovery—the ability to recognize when anomalous observations demand not just parameter updates but fundamental reconceptualization of the problem space itself.
Scientific discovery has always been a collaborative enterprise—across disciplines, institutions, and generations. AI systems represent new kinds of collaborative tools. The transition from static models to living systems marks a fundamental shift in how we conceive of AI systems that persist, that remember, that build intuition through repeated engagement with reality. Just as human scientists develop insight through years of experimentation, future AI systems will accumulate wisdom through continuous cycles of hypothesis, experiment, and revision. We thus call for the creation of new benchmarks and research programs centered around the proposed stacked architecture, moving evaluation beyond static datasets to interactive, discovery-oriented environments.
@@ 5 Outlook
Finally, it has to be emphasized that modern AI systems are already useful in their present form, and are being utilized effectively by scientific research groups across the world. However, even with future improvments, these tools bring many systemic hazards [41]: a) false positives and false negatives: spurious correlations can be mistaken for laws, while cautious priors may hide real effects, and thus rigorous uncertainty metrics and adversarial falsification must be built in; b) epistemic overconfidence: large models could shrink error bars off-distribution, demanding ensemble disagreement; c) erosion of insight and rigor: over time, there is signficiant risk of researchers losing key scientific skills; d) Cost: simulation-driven exploration can consume resources long after marginal information saturates, so schedulers must weigh value against resources; e) concept drift: equipments and sensors evolve, and thus without continual residual checks and rapid retraining, predictions may silently bias. These issues have to be continually acknowledged, recognizing and safeguards should be embeded into the scientific process.

### v3

@@ Abstract
The rapid evolution of artificial intelligence has led to expectations of transformative impact on science, yet current systems remain fundamentally limited in enabling genuine scientific discovery. This perspective contends that progress turns on closing three mutually reinforcing gaps in abstraction, reasoning and empirical grounding. Central to addressing these gaps is recognizing complementary cognitive modes: thinking as slow, iterative hypothesis generation—exploring counterfactual spaces where physical laws can be temporarily violated to discover new patterns—and reasoning as fast, deterministic validation, traversing established knowledge graphs to test consistency with known principles. Abstractions in this loop should be manipulable models that enable counterfactual prediction, causal attribution, and refinement. Design principles—rather than a monolithic recipe— are proposed for systems that reason in imaginary spaces and learn from the world: causal, multimodal models for internal simulation; persistent, uncertainty-aware scientific memory that distinguishes hypotheses from established claims; formal verification pathways coupled to computations and experiments. It is also argued that the inherent ambiguity in feedback from simulations and experiments, and underlying uncertainties make human judgment indispensable, not as a temporary scaffold but as a permanent architectural component. Evaluations must assess the system’s ability to identify novel phenomena, propose falsifiable hypotheses, and efficiently guide experimental programs toward genuine discoveries.
@@
Against this backdrop, this perspective piece is organized around three interlocking hurdles (i) the abstraction gap, which separates low-level statistical regularities from the mechanistic concepts on which scientists actually reason; (ii) the reasoning gap, which limits today’s models to correlation-driven pattern completion rather than causal, counterfactual inference; and (iii) the reality gap, which isolates computation from the empirical feedback loops that ultimately arbitrate truth. The central thesis is that scientific discovery demands a reimagining of AI architectures. Future AI systems must integrate active inference principles, maintaining persistent scientific memories while engaging in closed-loop interaction with both simulated and physical worlds. The following sections examine each gap in detail before presenting an integrated architecture that addresses these challenges holistically.
1 Fundamental Gaps in AI Models
The path from current AI capabilities to genuine scientific discovery is obstructed by interconnected barriers that reflect deep architectural limitations rather than scaling or engineering challenges. These gaps are not independent failures but symptoms of a unified problem: current AI systems lack the cognitive architecture necessary for scientific thinking. Understanding these gaps requires recognizing that they form a mutually reinforcing system of constraints: without rich abstractions there is little substrate for reasoning, and without tight coupling to reality even the most elegant abstractions may drift toward irrelevance.
1.1 The Abstraction Gap
@@
The gap between correlation and causation represents perhaps the most fundamental challenge in automated scientific discovery. While current models excel at finding statistical regularities, scientific understanding requires the ability to reason about interventions—to ask not just “what correlates with what?" but “what happens when we change this?". Pearl’s causal hierarchy [49] distinguishes three levels of cognitive ability: association (seeing), intervention (doing), and counterfactuals (imagining). Current AI systems operate primarily at the associative level, occasionally reaching intervention through experimental design. True scientific reasoning requires all three, particularly the counterfactual ability to imagine alternative scenarios that violate observed correlations. This connects directly to ethologist Konrad Lorenz’s insight –first tied to learning systems by Scholkopf [58]– that thinking is fundamentally about acting in imaginary spaces where we can violate the constraints of observed data. This mental experimentation—impossible in physical reality but accessible in the imagination—forms the basis of scientific law formation, as explained in Inset 1.
Critically, this mental experimentation must integrate diverse modalities of scientific data and representation. Scientific phenomena manifest as continuous vector/tensor fields over space-time (e.g., velocity-pressure fields, concentration gradients, electromagnetic fields) interleaved with discrete events (chemical reactions, phase transitions) and symbolic constructs (reaction mechanisms, theoretical frameworks). Effective AI systems for discovery must therefore maintain multi-modal embeddings that seamlessly translate between these representational levels—from raw sensor data to mathematical abstractions to causal hypotheses—enabling the system to reason simultaneously across observational patterns, physical mechanisms, and theoretical principles.
1.2 The Reasoning Gap
Future systems must balance the complementary modes of thinking and reasoning as first-class architectural principles. Thinking—or slow, iterative discovery of new patterns—demands (i) world-model agents that can explore counterfactual spaces through mental simulation [26]; (ii) curiosity-driven mechanisms that reward pattern novelty over immediate task performance; and (iii) patience parameters that prevent premature convergence. Reasoning—the fast, deterministic traversal of pattern graphs—demands (i) efficient knowledge graph architectures with learned traversal policies; (ii) neuro-symbolic stacks that maintain both continuous representations and discrete logical structures[40]; and (iii) caching mechanisms that transform expensive thinking outcomes into rapid reasoning primitives. The interplay between these modes mirrors how scientists alternate between exploratory experimentation (thinking) and theoretical derivation (reasoning) as referenced in Inset 1.
@@ 1.2 The Reasoning Gap
Scientific progress thrives on disciplined risk: venturing beyond received wisdom while remaining falsifiable. Current alignment protocols deliberately dampen exploratory behaviour, biasing models toward safe completion of well-trodden trajectories. Controlled speculation frameworks—for example, curiosity-driven reinforcement learning [46] combined with Bayesian epistemic uncertainty quantification—could allow systems to seek novel hypotheses, flag them with calibrated uncertainty, and propose targeted experiments for arbitration. Mechanisms such as self-consistency voting [64], adversarial peer review, and tool-augmented chain-of-thought audits offer additional scaffolding to keep high-variance reasoning connected to empirical reality.
A key aspect of closing the abstraction and reasoning gaps, therefore, is in developing architectures that can construct and manipulate explicit symbolic representations as dynamic objects rather than static patterns—essentially giving models the ability to retain abstract concepts like conservation laws or causal structures in working memory and actively transform them through mental experimentation [37]. This requires moving beyond current approaches that merely associate patterns to instead build compositional graph-based [5] reasoning systems where abstract principles can be instantiated, violated, and reconstructed in imaginary spaces. Future systems can thus target manipulable conceptual building blocks that can be assembled into novel configurations that have not been explicitly seen in training—enabling the kind of counterfactual reasoning that can ultimately bridge the gap between correlation and causation that defines scientific thinking.
1.3 The Reality Gap
While the abstraction and reasoning gaps constrain what AI systems can represent and manipulate internally, the reality gap addresses a more fundamental limitation: the disconnect between computational models and the real world they aim to describe. As detailed in Inset 2, the necessity of empirical feedback for scientific discovery emerges from theoretical constraints—Gödel’s incompleteness theorems and Wolfram’s computational irreducibility guarantee that no purely computational system can discover all truths about nature. Scientific progress depends not on escaping these constraints but on navigating them through continuous dialogue with reality. In practice, this dialogue is inherently multi-modal [18] and spatio-temporal [48, 27]. AI systems must maintain a physically grounded latent state, with dynamics constrained by known operators and invariances (conservation laws; symmetries such as Galilean/rotational). This structured, multi-modal view reduces sample complexity and yields representations that extrapolate across space, time, and interventions.
Empirical feedback complements formal reasoning by supplying information inaccessible to purely deductive systems, thereby expanding—rather than mechanically escaping—the set of testable scientific propositions. The interplay between formal systems and empirical validation creates a bootstrap mechanism that circumvents incompleteness and irreducibility constraints. This suggests that AI systems for discovery must be fundamentally open—not just to new data, but to surprise from reality itself. Scientific history abounds with internally coherent theories that later failed empirical tests, underscoring the indispensability of continuous validation against data. Current AI systems excel at interpolation within their training distributions but struggle with the extrapolation that defines discovery. This is exacerbated by the fact that many scientific domains are characterized by sparse, expensive data and imperfect simulators. Unlike language modeling where data is abundant, a single protein crystallography experiment might take months and cost thousands of dollars. Simulations help but introduce their own biases, further contributing to the reality gap.
The synthesis presented in Inset 2 directly informs our architecture. Thinking explores for new pockets and tests boundaries; reasoning exploits discovered regularities. World models encode provisional maps of known pockets, subject to Popper’s falsification and Kuhn’s paradigm shifts. Human steering proves essential. Humans provide non-computational insight for recognizing genuine understanding, value judgments for directing exploration, and navigation through paradigm shifts where evaluation criteria themselves transform. Humans can shape the search process by encoding domain knowledge, identifying significant anomalies, and recognizing connections that form larger frameworks. When Faraday discovered electromagnetic induction, he did not deduce it from Maxwell’s equations (which did not yet exist)—he found it through experiment. Thus, productive collaborations can implement the complete scientific method: AI generates and tests hypotheses at scale; humans provide insight and judgment and empirical feedback provides critical steering. Our architecture must therefore implement a hybrid loop: physics priors guide ML surrogates, which direct active experiments, which update our understanding in continuous iteration.
Physics priors
While generative models can create visually compelling outputs, they lack physical consistency—objects appear and disappear, gravity works intermittently, and causality is merely suggested rather than enforced. Mitchell [43] states that without biases to prefer some generalizations over others, a learning system cannot make the inductive leap necessary to classify instances beyond those it has already seen. Such inductive biases or physics priors—can be built-in to ensure generated realizations obey conservation laws, maintain object permanence, and support counterfactual reasoning about physical interactions.
Causal models
The current paradigm of domain-specific foundation models—from protein language models to molecular transformers—represents significant progress in encoding domain knowledge. However, these models fundamentally learn correlational patterns rather than causal mechanisms. ChemBERTa [14] can predict molecular properties through pattern matching but cannot simulate how modifying a functional group alters reaction pathways. AlphaFold [30] predicts protein structures through evolutionary patterns but does not model the physical folding process.
Scientific discovery requires models that transcend pattern recognition to capture causal dynamics. A causal molecular model would not just recognize that certain molecular structures correlate with properties—it would explain how electron density distributions cause reactivity, and how thermodynamic gradients drive reactions. This causal understanding enables the counterfactual reasoning essential to science: predicting outcomes of novel interventions never seen in training data. This architectural choice has profound implications: foundation models scale with data and compute, but causal models scale with understanding. With accumulaion of structural data, AI models can improve at interpolation. As we refine causal mechanisms, foundation models can improve at extrapolation—the essence of scientific discovery.
2 Additional Improvements
A certain level of consensus appears to be forming in the community that incremental scaling of present architectures may not deliver the qualitative leap that scientific discovery demands. Progress hinges on removing the design constraints through concurrent advances in algorithms and architectures (such as those described above) but also by improving efficiencies via hardware-software co-design and better evaluation benchmarks.
Computational Efficiency
@@ 2 Additional Improvements
The von Neumann bottleneck—shuttling tensors between distant memory and compute—now dominates energy budgets [41]. Processing-in-memory fabrics [34], spiking neuromorphic cores that exploit event sparsity, analog photonic accelerators for low-latency matrix products, quantum samplers for combinatorial sub-routines [2] could open new algorithmic spaces. Realising their potential outside of niche applications, however, will require co-design of hardware, software and algorithms and extensive community effort.
@@
3 Architecture of Active Inference AI Systems
The fundamental gaps analyzed in the preceding sections are not independent failures but symptoms of a deeper architectural mismatch between current AI systems and the requirements of scientific discovery. These insights, combined with the inherently multi-scale and multi-modal nature of scientific phenomena—from molecular interactions to emergent spatio-temporal dynamics—dictate specific architectural requirements. A system capable of genuine discovery must therefore integrate: internal models that support mental experimentation, knowledge structures that grow through thinking-reasoning cycles, and verification mechanisms that ground speculation in empirical reality. No monolithic approach can address these diverse demands; instead, we propose a modular active inference architecture where specialized components work in concert. Figure 1 illustrates this architecture, and the key components include:
@@ 3 Architecture of Active Inference AI Systems
The architectural principles outlined above find grounding in recent work on transformational scientific creativity. For instance, Schapiro et al. [56] formalize scientific conceptual spaces as directed acyclic graphs, where vertices represent generative rules and edges capture logical dependencies. This offers a concrete implementation pathway for the proposed dynamic knowledge graphs. Their distinction between modifying existing constraints versus fundamentally restructuring the space itself maps directly onto our architecture’s dual modes of reasoning (traversing established knowledge) and thinking (discovering new patterns that may violate existing assumptions). This convergence suggests that achieving transformational scientific discovery through AI systems requires systems capable of identifying and modifying the foundational axioms that constrain current scientific understanding—a capability the active inference framework aims to provide through its stacked architecture and integration of models, empirical feedback, and human guidance.
It is acknowledged that while the AI system can –in principle– be operated autonomously through well-defined interfaces between components, human interaction and decisions can be expected to play a key role. The architectural principles outlined above find partial instantiation in contemporary systems, though none fully realize the complete vision of scientific intelligence. Appendix A examines some current implementations through the lens of our three-gap framework, and discusses both substantial progress and persistent limitations that illuminate the path forward.
4 Limitations
@@
Finally, it has to be emphasized that modern AI systems are already useful in their present form, and are being utilized effectively by scientific research groups across the world. However, even with future improvements, these tools bring many systemic hazards [42]: a) false positives and false negatives: spurious correlations can be mistaken for laws, while cautious priors may hide real effects, and thus rigorous uncertainty metrics and adversarial falsification must be built in; b) epistemic overconfidence: large models could shrink error bars off-distribution, demanding ensemble disagreement; c) erosion of insight and rigor: over time, there is signficiant risk of researchers losing key scientific skills; d) Cost: simulation-driven exploration can consume resources long after marginal information saturates, so schedulers must weigh value against resources; e) concept drift: equipments and sensors evolve, and thus without continual residual checks and rapid retraining, predictions may silently bias. These issues have to be continually acknowledged, recognizing and safeguards should be embeded into the scientific process.
5 Summary and Outlook
This work has argued that the path to genuine scientific discovery through AI requires more than improvements in data, compute and scaling—it demands a reimagining of the underlying architecture. Current systems lack (i) abstractions that support mechanism-level reasoning, (ii) reasoning that operates counterfactually rather than correlationally, and (iii) reality coupling– particularly across multimodal, spatio-temporal measurements— that continuously calibrates beliefs to experiments and high-fidelity simulations. Although fully resolving these challenges is a long-horizon research program, immediate progress is possible by adopting an active-inference stack that (a) learns causal, multi-modal world models for internal simulation; (b) maintains persistent, versioned knowledge graphs with uncertainty; (c) routes claims either to formal proof engines or to empirical verification; and (d) operates under human steering where ambiguity and value trade-offs dominate. The aim is a shift from pattern completion to principle discovery.
This perspective calls upon the community to build on substantial progress in causal machine learning, active learning, and automated scientific discovery to address these critical gaps. The causal machine learning community has made significant strides in developing methods for causal inference from observational data, with frameworks such as Pearl’s causal hierarchy and recent advances in causal representation learning providing mathematical foundations for understanding interventions and counterfactuals. Similarly, active learning has evolved sophisticated strategies for optimal experimental design, while automated discovery systems have demonstrated success in specific domains such as materials science and drug discovery. However, these communities have largely operated in isolation, with causal methods focusing primarily on statistical inference rather than physical mechanism discovery, active learning optimizing for narrow uncertainty reduction rather than conceptual breakthroughs, and automated discovery systems excelling at interpolation within known spaces rather than extrapolation to genuinely novel phenomena. Realizing such a capability requires a diverse consortium to co-develop models, agentic infrastructure, formal verification stacks, simulators, robotic labs, and evaluations. Success cannot merely be measured by benchmarks alone, but by the moment when these systems become truly useful –as judged by domain experts– and make genuinely novel scientific discoveries.